Two charged particles are placed at a distance 1-0 cm apart. What is minimum possible magnitude of the electric force acting on each charge ?

Grade:12

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The minimum possible force would be given if the charges are 1 C.
So, the net force would be given by:
F=\frac{k\times 1^{2}}{r^{2}}
This gives us:
F=\frac{k\times 1^{2}}{r^{2}}=\frac{9\times 10^{9}}{\left (1\times 10^{-2} \right )^{2}}=9\times 10^{13} N
Regards
Sumit